### v1

# Working Backwards: Learning to Place by Picking

Abstract
We present Learning to Place by Picking (LPP), a method capable of autonomously collecting demonstrations for a family of placing tasks in which objects must be manipulated to specific locations. With LPP, we approach the learning of robotic object placement policies by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects that are initially located at their target placement locations. Our system is capable of collecting hundreds of demonstrations without human intervention by using a combination of tactile sensing and compliant control for grasps. We train a policy directly from visual observations through behaviour cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table and not at the original placement location). We validate our approach on home robotic scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
I Introduction
Object placing is an essential everyday task that autonomous robot manipulators will need to perform reliably. Many household chores, such as setting the table and loading a dishwasher, for example, require careful placing. Robotic placement involves contact-rich interactions and typically necessitates sensory feedback to determine the state of the in-hand object and the environment. Designing object placement policies is non-trivial and can involve solving a number of complicated subtasks, including object tracking, scene understanding, motion planning, and control.
Imitation learning (IL) provides an appealing, simple, and practical alternative by learning end-to-end control policies from expert demonstrations [1, 2]. The real world is complex and demonstration data should ideally include diverse examples to account for varying factors such as background, lighting conditions, and object types. Some recently-developed robotic systems [3, 4] have demonstrated the use of IL with a large amount of demonstration data to effectively solve a variety of tasks. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect expert demonstrations.
We propose a novel approach named Learning to Place by Picking (LPP) that facilitates and automates the collection of expert demonstrations for a large subset of placing problems by taking advantage of powerful grasp planners [5]. LABEL:fig:title provides a visual summary of our approach. LPP is a simple self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems. Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. In this way, we are able to generate place demonstration data in a self-supervised manner—the picking phase provides the training supervision for the place task. Finally, we learn a placing policy based only on visual observations using the demonstrations collected in the training environment. Crucially, this allows our reactive policy to be used for object placement outside of the training environment and in scenarios where the objects are not initially placed at their target locations (i.e., without privileged information). Our main contributions are:
  1. 1.
    LPP, a self-supervised data collection method for 6-DOF robotic object placements,
  2. 2.
    a flexible grasp planning pipeline that enables a user to specify what to grasp in a scene,
  3. 3.
    a tactile-driven regrasping and compliant control strategy to achieve uninterrupted pick and place data collection without human intervention, and
  4. 4.
    real-world experimental validation of LPP by using the collected demonstrations to train a vision-based policy capable of placing multiple plates in a dishrack and of setting a table.
II Related Work
In this section, we review previous approaches to object placing with manipulators and other work that generally leverage the idea of “working backwards”.
II-A Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [6, 7, 8, 9], while placing is relatively less explored.
Previous approaches have explored the use of motion planning combined with place-specific objectives [10, 11]. For example, in [10], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [12], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time. Other works have looked into combining learning with motion planning to operate directly on point clouds, thereby relaxing the assumption of known object and environment models [13, 14].
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [15] and imitation learning (IL) [16, 17]. In [16], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in combination with any existing methods that require expert demonstrations.
II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18, 19] and in learning visual predictive models for control [20, 21, 22]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate data. However, we focus on placement as opposed to insertion as our task. Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple objects and place poses to be involved in the task. Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
III Learning to Place by Picking
We first present an overview of our self-supervised data collection method, LPP, in Section III-A and describe a method for noise augmentation during data collection in Section III-B. We then provide details on the procedure used to train our IL policy in Section III-C.
III-A Self-supervised Data Collection
[…]
Figure 1: The four steps involved in LPP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [5], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
LPP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving and (4) placing by reversing; these steps are visualized in Figure 1. During training, the objects of interest are initially in their goal positions.
III-A1 Grasp Planning
[…]
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (a) Grounding-Dino [27] for object bounding box detection based on user provided texts and (b) Segment Anything [28] on the cropped images to produce object specific masks. We then use (c) Contact-graspnet [5] for grasp generation on only the segmented objects. Top: We specify “green plate” and “blue plate” as the targets for data collection. Bottom: We specify “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1. We use the grasp planner Contact-GraspNet [5] to generate $L$ candidate grasp poses $\mathbf{T}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$. The user then provides text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”). Given the user-provided texts, we use Grounding-Dino [27] to find the object bounding boxes. Finally, we run Segment Anything [28] on each of the previously generated image crops from the bounding boxes to generate per-object masks, which Contact-GraspNet can then use to filter grasps that are not of objects of interest: $\bar{\mathbf{T}}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$ and $\bar{\mathbf{T}}_{g}\subseteq\mathbf{T}_{g}$. We visualize an example of this process in Figure 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{g,k}$, from $\bar{\mathbf{T}}_{g}$. To approach the object of interest without collisions, we include a pregrasp pose, $\mathbf{T}^{\text{pre}}_{g,k}$, that is defined with a translational offset above the grasp pose $\mathbf{T}_{g,k}$.
III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Figure 1. During grasping, two modules play a critical role in the robustness of LPP: compliant control for grasping (CCG) and tactile regrasping (TR).
Compliant Control for Grasping We use a cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. The low compliance of the robot arm then naturally allows the gripper to comply with the contact constraints and minimize its applied forces against the environment.
Tactile Regrasping Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip once grasped. To mitigate this issue, we use tactile sensing to preemptively detect if a grasp is stable and perform a regrasp if needed.
In Section IV-A, we experimentally investigate the effects of including these modules through ablation studies.
III-A3 Retrieving
We begin the retrieval phase when a valid stable grasp is detected. An image of the retrieval procedure is displayed in Figure 1. We store the pose of the end effector (EE) as well as the respective timestamp throughout the retrieval process. The manipulator is set to be compliant along the rotational axes of the EE when commanded to move to the pregrasp pose. Doing so allows the manipulator to better comply to any unexpected interactions with the environment. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{c,k}$ centered around a fixed pose above the scene. That is, given a fixed clearance pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$. We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection. At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
$\tau^{\text{expert}}_{r}=((\mathbf{T}_{0},t_{0}),...,(\mathbf{T}_{M},t_{M})),$ (1)
where $\mathbf{T}_{0}=\mathbf{T}_{g,k}$, $\mathbf{T}_{M}=\mathbf{T}_{c,k}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a more sparse trajectory $\bar{\tau}^{\text{expert}}_{r}$ of length $\bar{M}<M$. Our visual policies operate at 5 $\mathrm{h}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
III-A4 Placing by reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory:
$\bar{\tau}^{\text{expert}}_{p}=(\mathbf{T}_{\Delta t_{\bar{M}}},...,\mathbf{T}_{\Delta t_{0}}).$ (2)
An example of a place trajectory is shown in Figure 1. Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
$\mathbf{T}_{\Delta\text{cmd},i}=\mathbf{T}^{-1}_{\Delta t_{\bar{M}-i}}\mathbf{T}_{\Delta t_{\bar{M}-i-1}},$ (3)
where $i\in\{1,...,\bar{M}\}$ and $\mathbf{T}_{\Delta\text{cmd},i}\in SE(3)$ is a relative change in EE pose. In practice, we use six-dimensional vectors $[\mathbf{t},\boldsymbol{\theta}]^{T}$ to represent poses, where $\mathbf{t}\in\mathbb{R}^{3}$ is a translation vector and $\boldsymbol{\theta}\in\mathbb{R}^{3}$ is a rotation vector. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}^{1}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We define the expert action as $\mathbf{a}_{i}=(\mathbf{a}_{\Delta\text{cmd},i},a_{\text{gripper},i})$ where $\mathbf{a}_{\Delta\text{cmd},i}=[\mathbf{t}_{i},\boldsymbol{\theta}_{i}]^{T}$ is the six-dimensional vector representation of the pose $\mathbf{T}_{\Delta\text{cmd},i}$ and $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}}$, and store the sequences of observations and actions as an expert trajectory. We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$, where $\mathbf{o}_{i}\in\mathbb{R}^{w\times h\times 3}$. We downsample the images to a resolution of $128\times 128$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper for N time steps (i.e. $\mathbf{a}_{\Delta\text{cmd},j}=[0,0,0,0,0,0]^{T}$ and $a_{\text{gripper},j}=0$ for all $j\in\{\bar{M}+1,...,\bar{M}+N\}$). We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples:
$\tau^{expert}=\{\mathbf{o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1}\}_{t=0}^{T},$ (4)
with a total length of $T=\bar{M}+N$.
III-B Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise. We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [29, 30], and actions or controls [31]. We call this variant noise-augmented data collection. We perturb the translation of the pose $\mathbf{t}$ as follows:
$\mathbf{t}_{perturbed}=\mathbf{t}+\delta\mathbf{t},$ (5)
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation as follows:
$\boldsymbol{\theta}_{perturbed}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (6)
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We find $\sigma_{t}=0.5$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-C Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}|\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by LPP. We use behavioral cloning (BC) [32] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}|\mathbf{o}_{t})}.$ (7)
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [33], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [34, 35, 36], by setting the standard deviation of the Gaussian distributions to be the arbitrarily small value of $61e-$ 6. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss. As commonly done [37], we stack a history of RGB frames into sequences of length four for our observations.
IV Experiments
[…]
Figure 3: Sequence of images from roll outs of place policies trained using data collected with LPP. The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
We use LPP to train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 3. In both cases, the place policy has to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we present our quantitative results for the dishrack loading task as it better represents the type of task, with contact constraints from the environment, that LPP was designed for. We study the robustness of LPP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the quality the policies trained with data from LPP and kinesthetic teaching in Section IV-C.
IV-A LPP Data Collection Robustness
Figure 4: Tactile images from before and after a tactile regrasp (TR). Left: the plate’s edge is seen in the middle of the tactile image indicating a shallow grasp. Right: after a regrasp, the edge of the plate has shifted to a desired spot indicating a stable grasp.
Figure 5: Ablation study on compliant control for grasping (CCG) and tactile regrasp (TR). We measure the average number of episodes that can be collected autonomously based on the number of failures. Both CCG and TR play a crucial role in reducing the number of failures to zero, which allows us to collect the necessary amount of data seamlessly.
We require a robust data collection loop to collect a large amount of place demonstrations autonomously with LPP. We study the effects of compliant control for grasping (CCG) and tactile regrasp (TR), as introduced in Section III, on the robustness of LPP.
We observed two main causes that led to misplaced plates during data collection. First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot. With LPP, we mitigate this failure mode with CCG where we decrease the translational and rotational stiffness of the manipulator controller right before grasping the object. The second failure mode involved unstable grasps that lead to large relative motion between the object and the gripper. With LPP, we mitigate unstable grasps using TR to preemptively detect if a grasp will lead to relative object motion and to regrasp if needed. Based on empirical data, we find that unstable grasps typically occurred with shallow grasps. To this end, we design our regrasping module to recognize edges and calculate the relative change in EE pose required to recover a deep enough stable grasp.
[…]
Figure 6: Visualization of edges in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
To determine the position of an edge, we fit an ellipse to the detected contact region. If an edge is detected, we calculate the difference between the ideal and detected depth of the edge and use this to calculate the commanded change in EE pose. An example of this is shown in Figure 4. While not all graspable objects contain edges, a large subset of items do as shown in Figure 6.
We conduct an ablation study on CCG and TR where we run self-supervised data collection for a total of 128 episodes and record the number of failures. We plot the results in Figure 5. The naive approach without CCG and without TR fails a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The amount of failures drop to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention. Finally, with both CCG and TR, we are able to successfully collect 128 episodes without any human intervention.
IV-B Noise Augmentation Ablation
TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation using two different policy action representations. The noise-augmented data collection procedure improves the performance of both deterministic and Gaussian mixture policies.
Action Representation Trained with Noise Aug. Data Avg. Success Rate ($\uparrow$) Total Successes ($\uparrow$)
Deterministic No 71.67(5.93) 57/80
Yes 81.11(5.50) 56/70
Gaussian Mixture No 72.22(5.67) 50/70
Yes 87.78(3.93) 62/70
[…]
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [38]. We also qualitatively observe that policies trained with noise-augmented datasets tend to be better at re-adjusting the object with respect to the goal.
IV-C Comparison to Kinesthetic Teaching
Figure 7: Comparison of success rates of policies trained using a dataset collected by LPP and a traditional kinesthetic teaching approach. We test models trained on an increasing amount of demonstration data. Policies trained on data collected by LPP outperform those trained on data collected from kinesthetic teaching in almost all dataset sizes. The shaded region consists of one standard deviation.
We compare the performance of policies trained using demonstration data from LPP and from kinesthetic teaching. In particular, we investigate whether the quality of the demonstrations differ. To compare, we train two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with LPP and another collected with kinesthetic teaching [39]. We use a Gaussian mixture model as the action representation of both policies. For LPP, we use the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
[…]
We hypothesize that the kinesthetic policy performs more poorly than the LPP policy because the overall quality of the human demonstrations is lower than the programmatic or machine-generated demonstrations produced by LPP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [40, 41, 42]. In particular, human demonstrations can be sub-optimal due to human mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think).
V Limitations and Future Work
The quality and diversity of the collected grasps in LPP are partly determined by the viewpoint of the camera. In this work, we empirically select a viewpoint that is adequate for our use-case. It would be interesting to allow the manipulator to explore and find vantage points for better grasp generation. Furthermore, while tactile sensing is used to improved the robustness of data collection, it could also be used as an extra modality for the policy.
VI Conclusion
We presented LPP, a self-supervised data collection method to collect expert demonstrations for placement tasks. LPP leverages recent advancement in grasp planners and the symmetry between picking and placing to autonomously collect placing demonstrations. LPP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by LPP surpasses that of human demonstrations collected by kinesthetic teaching. LPP provides a promising and simple direction for real world robotic placing demonstration data collection with minimal human labour.

### v2

@@
 Abstract
-We present Learning to Place by Picking (LPP), a method capable of autonomously collecting demonstrations for a family of placing tasks in which objects must be manipulated to specific locations.
-With LPP, we approach the learning of robotic object placement policies by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems.
-Specifically, we obtain placing demonstrations from a set of grasp sequences of objects that are initially located at their target placement locations.
-Our system is capable of collecting hundreds of demonstrations without human intervention by using a combination of tactile sensing and compliant control for grasps.
-We train a policy directly from visual observations through behaviour cloning, using the autonomously-collected demonstrations.
-By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table and not at the original placement location).
-We validate our approach on home robotic scenarios that include dishwasher loading and table setting.
+We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific contact-constrained locations.
+With PvP, we approach the collection of robotic object placement demonstrations by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems.
+Specifically, we obtain placing demonstrations from a set of grasp sequences of objects initially located at their target placement locations.
+Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention by combining two modules: tactile regrasping and compliant control for grasps.
+We train a policy directly from visual observations through behavioral cloning, using the autonomously-collected demonstrations.
+By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table).
+We validate our approach in home robotic scenarios that include dishwasher loading and table setting.
 Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
 I Introduction
-Object placing is an essential everyday task that autonomous robot manipulators will need to perform reliably.
-Many household chores, such as setting the table and loading a dishwasher, for example, require careful placing.
-Robotic placement involves contact-rich interactions and typically necessitates sensory feedback to determine the state of the in-hand object and the environment.
-Designing object placement policies is non-trivial and can involve solving a number of complicated subtasks, including object tracking, scene understanding, motion planning, and control.
-Imitation learning (IL) provides an appealing, simple, and practical alternative by learning end-to-end control policies from expert demonstrations [1, 2].
-The real world is complex and demonstration data should ideally include diverse examples to account for varying factors such as background, lighting conditions, and object types.
-Some recently-developed robotic systems [3, 4] have demonstrated the use of IL with a large amount of demonstration data to effectively solve a variety of tasks.
-A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect expert demonstrations.
-We propose a novel approach named Learning to Place by Picking (LPP) that facilitates and automates the collection of expert demonstrations for a large subset of placing problems by taking advantage of powerful grasp planners [5].
+Imitation learning (IL) provides an appealing, simple, and practical option by learning end-to-end control policies from expert demonstrations [1, 2].
+The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics.
+While there have been many attempts to transfer this paradigm to robotic systems [3, 4], robots have yet to achieve similar successes.
+A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect the large amount of expert demonstrations in the physical world.
+Having said this, many recent methods attempt to streamline [5, 6] or automate [7, 8, 9, 10] data collection.
+As highlighted by the authors of [9], the lack of robust and diverse autonomous data collection policies is a major limitation to scaling up IL.
+In this work, we are interested in investigating autonomous data collection for robotic object placement.
+As opposed to prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints.
+We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place if handled properly.
+We propose a novel approach named placing via picking (PvP) that automates the collection of expert demonstrations for a large subset of contact-constrained placing problems.
+We do so by taking advantage of a powerful grasp planner [11], tactile sensing, and compliant control.
+PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems.
+Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory.
+We generate demonstration data for object placement in a self-supervised manner—the picking phase provides the training supervision for the placing task.
+While appearing deceptively easy at first, we highlight the crucial importance of two modules, compliant control for grasping (CCG) and tactile regrasping (TR).
+CCG and TR enables (1) robust and uninterrupted pick-and-place and (2) leveraging environment contact-constraints as guides to inform object placement.
 LABEL:fig:title provides a visual summary of our approach.
-LPP is a simple self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems.
-Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory.
-In this way, we are able to generate place demonstration data in a self-supervised manner—the picking phase provides the training supervision for the place task.
-Finally, we learn a placing policy based only on visual observations using the demonstrations collected in the training environment.
-Crucially, this allows our reactive policy to be used for object placement outside of the training environment and in scenarios where the objects are not initially placed at their target locations (i.e., without privileged information).
 Our main contributions are:
   1. 1.
-    LPP, a self-supervised data collection method for 6-DOF robotic object placements,
+    PvP, a self-supervised data collection method for 6-DOF robotic object placements;
   2. 2.
-    a flexible grasp planning pipeline that enables a user to specify what to grasp in a scene,
+    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection without human intervention in contact-constrained environments;
   3. 3.
-    a tactile-driven regrasping and compliant control strategy to achieve uninterrupted pick and place data collection without human intervention, and
+    a language-specified grasp planning pipeline; and
   4. 4.
-    real-world experimental validation of LPP by using the collected demonstrations to train a vision-based policy capable of placing multiple plates in a dishrack and of setting a table.
+    real-world experimental validation of PvP using collected demonstrations to train a vision-based policy capable of placing multiple plates in a dish rack and of setting a table.
 II Related Work
-In this section, we review previous approaches to object placing with manipulators and other work that generally leverage the idea of “working backwards”.
-II-A Robotic Manipulator Placing
-Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [6, 7, 8, 9], while placing is relatively less explored.
+In this section, we review previous approaches to automate robot data collection in the real world and other work that generally leverage the idea of “working backwards”.
+II-A Automatic Data Collection
+A large number of existing policy learning approaches use prior policies to automate the data collection process.
+Examples include using trajectory optimization [12, 13] and, closer to our work, simple scripted pick-and-place policies [7, 8, 9] as forms of supervision.
+The authors of [7] use a scripted pick-and-place policy to bootstrap the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator.
+In [9], the authors present AutoRT, a method that uses a combination of scripted, learned and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data.
+Using a combination of a large language model (LLM) and visual language models (VLMs), AutoRT observes scenes in the wild and comes up with plausible tasks.
+Once a task is defined and chosen, AutoRT samples a data collection policy to attempt the task and collect real world data.
+In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with potential contact constraints.
+Unlike prior work, we do not restrict the placing task to simpler small objects on flat horizontal surfaces.
+The authors of [9] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; Our work is complementary with many of the prior frameworks and PvP could be used as a prior policy within them.
+II-B Working Backwards
+The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [14] and in learning visual predictive models for control [15, 16, 17].
+Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [18, 19, 20].
+These methods consider objects that are assumed to be in their desired place configuration during data collection.
+The manipulator can then pick an object and move back to the initial goal location to generate data.
+However, we focus on placement as opposed to insertion as our task.
+Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple, generally larger objects and place poses involved in the task.
+Furthermore, unlike prior work [18, 19, 20], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning.
+Our approach also resembles [21], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner.
+However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement.
+In contrast, the self-supervised data collection procedure introduced in [21] uses a 2D open loop pick-and-place action primitive.
+II-C Robotic Manipulator Placing
+Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [22, 23, 24], while placing is relatively less explored.
 Previous approaches have explored the use of motion planning combined with place-specific objectives [25, 26].
@@
 Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time.
-Other works have looked into combining learning with motion planning to operate directly on point clouds, thereby relaxing the assumption of known object and environment models [13, 14].
-Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [15] and imitation learning (IL) [16, 17].
+Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [28] and imitation learning (IL) [29].
 In [29], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing.
-The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in combination with any existing methods that require expert demonstrations.
-II-B Working Backwards
-The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18, 19] and in learning visual predictive models for control [20, 21, 22].
-Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25].
-These methods consider objects that are assumed to be in their desired place configuration during data collection.
-The manipulator can then pick an object and move back to the initial goal location to generate data.
-However, we focus on placement as opposed to insertion as our task.
-Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple objects and place poses to be involved in the task.
-Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning.
-Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner.
-However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement.
-In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
-III Learning to Place by Picking
-We first present an overview of our self-supervised data collection method, LPP, in Section III-A and describe a method for noise augmentation during data collection in Section III-B.
-We then provide details on the procedure used to train our IL policy in Section III-C.
+The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with any methods that require expert demonstrations.
+III Placing via Picking
+We first present an overview of our self-supervised data collection method, PvP, in Section III-A and then describe a method for noise augmentation during data collection in Section III-D.
+Next, we provide details about the procedure used to train our IL policy in Section III-E.
 III-A Self-supervised Data Collection
@@
 […]
-Figure 1: The four steps involved in LPP, our autonomous demonstration data collection process for placing.
+Figure 1: The four steps involved in PvP, our autonomous demonstration data collection process for placing.
 We (a) generate grasps with an off-the-shelf grasp planner [11], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
-LPP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving and (4) placing by reversing; these steps are visualized in Figure 1.
+PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Figure 1.
 During training, the objects of interest are initially in their goal positions.
@@
 Figure 2: The steps involved in our language-driven grasp planning pipeline.
-We use (a) Grounding-Dino [27] for object bounding box detection based on user provided texts and (b) Segment Anything [28] on the cropped images to produce object specific masks.
-We then use (c) Contact-graspnet [5] for grasp generation on only the segmented objects.
-Top: We specify “green plate” and “blue plate” as the targets for data collection.
-Bottom: We specify “bowl” and “mug” as the targets for data collection.
+We use (a) Grounding-Dino [30] for object bounding box detection based on text and (b) Segment Anything [31] on the cropped images to produce object specific masks.
+We then use (c) Contact-graspnet [11] for grasp generation on only the segmented objects (i.e., masked areas).
+Top: using “green plate” and “blue plate” as the targets for data collection.
+Bottom: using “bowl” and “mug” as the targets for data collection.
 At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1.
-We use the grasp planner Contact-GraspNet [5] to generate $L$ candidate grasp poses $\mathbf{T}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$.
-The user then provides text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”).
-Given the user-provided texts, we use Grounding-Dino [27] to find the object bounding boxes.
-Finally, we run Segment Anything [28] on each of the previously generated image crops from the bounding boxes to generate per-object masks, which Contact-GraspNet can then use to filter grasps that are not of objects of interest: $\bar{\mathbf{T}}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$ and $\bar{\mathbf{T}}_{g}\subseteq\mathbf{T}_{g}$.
+We use the grasp planner Contact-GraspNet [11] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$.
+Given a set of text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [30] to find the object bounding boxes.
+Finally, we run Segment Anything [31] on each of the previously generated image crops from the bounding boxes to generate per-object masks.
+Contact-GraspNet can then use these masks to filter grasps that are not on objects of interest and keep only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$.
 We visualize an example of this process in Figure 2 for two different scenes.
-After pruning, we randomly select a single grasp pose, $\mathbf{T}_{g,k}$, from $\bar{\mathbf{T}}_{g}$.
-To approach the object of interest without collisions, we include a pregrasp pose, $\mathbf{T}^{\text{pre}}_{g,k}$, that is defined with a translational offset above the grasp pose $\mathbf{T}_{g,k}$.
+After pruning, we randomly select a single grasp pose, $\mathbf{T}_{\text{grasp}}$, from $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$.
+To approach the object of interest without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
 III-A2 Grasping
@@
 An example of a grasp sequence is shown in Figure 1.
-During grasping, two modules play a critical role in the robustness of LPP: compliant control for grasping (CCG) and tactile regrasping (TR).
-Compliant Control for Grasping We use a cartesian impedance controller for manipulation.
+During grasping, two modules play a critical role in the robustness of PvP: compliant control for grasping (CCG) and tactile regrasping (TR).
+III-B Compliant Control for Grasping
+We use a cartesian impedance controller for manipulation.
 This choice allows us to set the level of compliance of the manipulator at various stages as needed.
 Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values.
-The low compliance of the robot arm then naturally allows the gripper to comply with the contact constraints and minimize its applied forces against the environment.
-Tactile Regrasping Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip once grasped.
-To mitigate this issue, we use tactile sensing to preemptively detect if a grasp is stable and perform a regrasp if needed.
+The low compliance of the robot arm then allows the manipulator and gripper to comply with the contact constraints and minimize its applied forces against the environment.
+Larger contact-constrained objects generate unnecessary reaction forces when gripped without compliance.
+These large forces trigger emergency stops and can damage the environment, and the robot.
+Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper ultimately leading to inaccurate placements.
+On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator and gripper towards a natural object placement pose.
+Figure 3: Tactile images from before and after a tactile regrasp (TR).
+Left: the plate’s contact surface area fills half of the tactile image indicating a shallow grasp.
+Right: after a regrasp, the contact surface area has increased indicating a stable grasp.
+[…]
+[…]
+[…]
+[…]
+Figure 4: Visualization of contact surfaces in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
+III-C Tactile Regrasping
+Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip significantly once grasped.
+In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) results in grasps that are unstable (e.g., a shallow grasp) and not ideal for PvP.
+To mitigate this issue, we use tactile sensing (specifically, a visuotactile Finger-STS [32] sensor) to preemptively detect if a grasp is stable and perform a regrasp if needed.
+The Finger-STS is capable of providing multimodal feedback (i.e., both tactile and visual observations).
+To this end, we design our regrasping module to detect the contact surface area and calculate the relative change in EE pose required to recover a larger contact surface and, thereby, a more stable grasp.
+We detect regions of contact based on marker displacements and the RGB image from the Finger-STS.
+If an impartial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose.
+We note that a simpler and more general random regrasp strategy could also be used.
+An example of this is shown in Figure 3.
+We visualize contact surfaces of various household items in Figure 4.
 In Section IV-A, we experimentally investigate the effects of including these modules through ablation studies.
-III-A3 Retrieving
-We begin the retrieval phase when a valid stable grasp is detected.
-An image of the retrieval procedure is displayed in Figure 1.
-We store the pose of the end effector (EE) as well as the respective timestamp throughout the retrieval process.
-The manipulator is set to be compliant along the rotational axes of the EE when commanded to move to the pregrasp pose.
-Doing so allows the manipulator to better comply to any unexpected interactions with the environment.
-Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{c,k}$ centered around a fixed pose above the scene.
-That is, given a fixed clearance pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$.
+III-C1 Retrieving
+We begin the retrieval phase when a valid, stable grasp is detected.
+An image of the retrieval procedure is shown in Figure 1.
+We store the poses of the end effector (EE) as well as the respective timestamps throughout the retrieval process.
+For our work, we measure joint encoder readings at a rate of 120 $\mathrm{H}\mathrm{z}$.
+The manipulator is set to be compliant along the rotational axes of the EE only, when commanded to move to the pregrasp pose.
+This allows the manipulator to better comply with any unexpected environment interactions.
+Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ centered around a set fixed pose above the scene.
+That is, given a set fixed pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$ to recover a clearance pose.
 We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection.
@@
 $\tau^{\text{expert}}_{r}=((\mathbf{T}_{0},t_{0}),...,(\mathbf{T}_{M},t_{M})),$ (1)
-where $\mathbf{T}_{0}=\mathbf{T}_{g,k}$, $\mathbf{T}_{M}=\mathbf{T}_{c,k}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0.
-To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a more sparse trajectory $\bar{\tau}^{\text{expert}}_{r}$ of length $\bar{M}<M$.
-Our visual policies operate at 5 $\mathrm{h}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
-III-A4 Placing by reversing
-We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory:
-$\bar{\tau}^{\text{expert}}_{p}=(\mathbf{T}_{\Delta t_{\bar{M}}},...,\mathbf{T}_{\Delta t_{0}}).$ (2)
+where $\mathbf{T}_{0}=\mathbf{T}_{\text{grasp}}$, $\mathbf{T}_{M}=\mathbf{T}_{\text{clearance}}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0.
+To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a sparser trajectory $\bar{\tau}^{\text{expert}}_{r}$.
+We extract a total of $\bar{M}+1$ states where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$.
+We do this by finding the nearest pose in terms of time from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
+$\bar{\tau}^{\text{expert}}_{r}=((\bar{\mathbf{T}}_{0},0),(\bar{\mathbf{T}}_{1},\Delta t),...,(\bar{\mathbf{T}}_{\bar{M}},\bar{M}\Delta t)).$ (2)
+In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
+III-C2 Placing by reversing
+We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory
+$\bar{\tau}^{\text{expert}}_{p}=(\bar{\mathbf{T}}_{\bar{M}},...,\bar{\mathbf{T}}_{0}).$ (3)
 An example of a place trajectory is shown in Figure 1. Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
-$\mathbf{T}_{\Delta\text{cmd},i}=\mathbf{T}^{-1}_{\Delta t_{\bar{M}-i}}\mathbf{T}_{\Delta t_{\bar{M}-i-1}},$ (3)
-where $i\in\{1,...,\bar{M}\}$ and $\mathbf{T}_{\Delta\text{cmd},i}\in SE(3)$ is a relative change in EE pose.
-In practice, we use six-dimensional vectors $[\mathbf{t},\boldsymbol{\theta}]^{T}$ to represent poses, where $\mathbf{t}\in\mathbb{R}^{3}$ is a translation vector and $\boldsymbol{\theta}\in\mathbb{R}^{3}$ is a rotation vector.
-The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}^{1}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis.
-We define the expert action as $\mathbf{a}_{i}=(\mathbf{a}_{\Delta\text{cmd},i},a_{\text{gripper},i})$ where $\mathbf{a}_{\Delta\text{cmd},i}=[\mathbf{t}_{i},\boldsymbol{\theta}_{i}]^{T}$ is the six-dimensional vector representation of the pose $\mathbf{T}_{\Delta\text{cmd},i}$ and $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close).
-To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}}$, and store the sequences of observations and actions as an expert trajectory.
-We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$, where $\mathbf{o}_{i}\in\mathbb{R}^{w\times h\times 3}$.
-We downsample the images to a resolution of $128\times 128$.
-During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}\}$).
-Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper for N time steps (i.e.
-$\mathbf{a}_{\Delta\text{cmd},j}=[0,0,0,0,0,0]^{T}$ and $a_{\text{gripper},j}=0$ for all $j\in\{\bar{M}+1,...,\bar{M}+N\}$).
+$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}_{\bar{M}-i-1},$ (4)
+where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in SE(3)$ is a relative change in EE pose.
+We define the expert action as $\mathbf{a}_{i}=(\Delta\mathbf{T}_{i},a_{\text{gripper},i})$, where $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close).
+To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}-1}$, and store the sequences of observations and actions as an expert trajectory.
+We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$.
+We downsample the images to a resolution of $128\times 128$ and also stack the three previous RGB frames, which results in $\mathbf{o}_{i}\in\mathbb{R}^{128\times 128\times 12}$.
+During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$).
+Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper (i.e.
+$\Delta\mathbf{T}_{i}=\boldsymbol{\mathbf{I}}$ and $a_{\text{gripper},i}=0$) for N time steps.
 We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object.
-A single expert demonstration or trajectory is then defined as a set of training tuples:
-$\tau^{expert}=\{\mathbf{o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1}\}_{t=0}^{T},$ (4)
+A single expert demonstration or trajectory is then defined as a set of training tuples
+$\tau^{\text{expert}}=\{\mathbf{(o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1})\}_{t=0}^{T},$ (5)
 with a total length of $T=\bar{M}+N$.
-III-B Noise-Augmented Data Collection
+III-D Noise-Augmented Data Collection
 To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise.
-We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [29, 30], and actions or controls [31].
+We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [33], and actions [34].
+We also draw inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [35], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment.
 We call this variant noise-augmented data collection.
-We perturb the translation of the pose $\mathbf{t}$ as follows:
-$\mathbf{t}_{perturbed}=\mathbf{t}+\delta\mathbf{t},$ (5)
-where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$.
-Similarly, we perturb the rotation as follows:
-$\boldsymbol{\theta}_{perturbed}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (6)
-where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$.
-We find $\sigma_{t}=0.5$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values.
+In practice, we represent pose commands $\Delta\mathbf{T}$ as a combination of a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a rotation vector $\boldsymbol{\theta}\in\mathbb{R}^{3}$.
+The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis.
+We perturb the translation of the pose with
+$\mathbf{t}_{\text{perturbed}}=\mathbf{t}+\delta\mathbf{t},$ (6)
+where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation with
+$\boldsymbol{\theta}_{\text{perturbed}}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (7)
+where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$.
+We perturb both the rotation axis and the rotation angle.
+We find $\sigma_{t}=0.5~{}$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5~{}$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values.
 We evaluate the effects of this augmentation on policy performance in Section IV-B
-III-C Policy Learning
-We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}|\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by LPP.
+III-E Policy Learning
+We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP.
 We use behavioral cloning (BC) [36] to train our policy with a likelihood-based loss or objective function.
@@
 In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians.
-In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [34, 35, 36], by setting the standard deviation of the Gaussian distributions to be the arbitrarily small value of $61e-$ 6.
+In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [38] and [39], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value of $61e-$ 6.
 Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss.
-As commonly done [37], we stack a history of RGB frames into sequences of length four for our observations.
 IV Experiments
@@
 […]
-Figure 3: Sequence of images from roll outs of place policies trained using data collected with LPP.
+Figure 5: Sequence of images from roll outs of place policies trained using data collected with PvP.
 The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
-We use LPP to train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 3.
-In both cases, the place policy has to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle).
-In this section, we present our quantitative results for the dishrack loading task as it better represents the type of task, with contact constraints from the environment, that LPP was designed for.
-We study the robustness of LPP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the quality the policies trained with data from LPP and kinesthetic teaching in Section IV-C.
-IV-A LPP Data Collection Robustness
-Figure 4: Tactile images from before and after a tactile regrasp (TR).
-Left: the plate’s edge is seen in the middle of the tactile image indicating a shallow grasp.
-Right: after a regrasp, the edge of the plate has shifted to a desired spot indicating a stable grasp.
-Figure 5: Ablation study on compliant control for grasping (CCG) and tactile regrasp (TR).
-We measure the average number of episodes that can be collected autonomously based on the number of failures.
-Both CCG and TR play a crucial role in reducing the number of failures to zero, which allows us to collect the necessary amount of data seamlessly.
-We require a robust data collection loop to collect a large amount of place demonstrations autonomously with LPP.
-We study the effects of compliant control for grasping (CCG) and tactile regrasp (TR), as introduced in Section III, on the robustness of LPP.
-We observed two main causes that led to misplaced plates during data collection.
+We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 5.
+In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle).
+In this section, we present our quantitative results for the more challenging dishrack loading task as it better represents the type of task, with additional contact constraints from the environment, that PvP was designed for.
+We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
+IV-A PvP Data Collection Robustness
+Figure 6: Ablation study on compliant control for grasping (CCG) and tactile regrasping (TR).
+We measured the average number of episodes that could be collected autonomously based on the number of failures.
+Both CCG and TR play a crucial role in reducing the number of failures to zero, which allowed us to collect the necessary amount of data seamlessly.
+We require a robust data collection loop to collect a large number of place demonstrations autonomously with PvP.
+We studied the effects of compliant control for grasping (CCG) and tactile regrasping (TR), as introduced in Section III, on the robustness of PvP.
+We observed two main causes, initially discussed in Section III-A2, that led to misplaced plates during data collection.
 First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack.
@@
 This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot.
-With LPP, we mitigate this failure mode with CCG where we decrease the translational and rotational stiffness of the manipulator controller right before grasping the object.
-The second failure mode involved unstable grasps that lead to large relative motion between the object and the gripper.
-With LPP, we mitigate unstable grasps using TR to preemptively detect if a grasp will lead to relative object motion and to regrasp if needed.
-Based on empirical data, we find that unstable grasps typically occurred with shallow grasps.
-To this end, we design our regrasping module to recognize edges and calculate the relative change in EE pose required to recover a deep enough stable grasp.
-[…]
-[…]
-[…]
-[…]
-Figure 6: Visualization of edges in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
-To determine the position of an edge, we fit an ellipse to the detected contact region.
-If an edge is detected, we calculate the difference between the ideal and detected depth of the edge and use this to calculate the commanded change in EE pose.
-An example of this is shown in Figure 4.
-While not all graspable objects contain edges, a large subset of items do as shown in Figure 6.
-We conduct an ablation study on CCG and TR where we run self-supervised data collection for a total of 128 episodes and record the number of failures.
-We plot the results in Figure 5.
-The naive approach without CCG and without TR fails a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed.
-The amount of failures drop to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention.
-Finally, with both CCG and TR, we are able to successfully collect 128 episodes without any human intervention.
+With PvP, we mitigate this failure mode with CCG.
+The second failure mode involved unstable grasps that lead to a large relative motion between the object and the gripper.
+With PvP, we mitigate unstable grasps using TR to preemptively detect unstable grasps and to regrasp if needed.
+We conducted an ablation study on CCG and TR where we ran self-supervised data collection for a total of 128 episodes and recorded the number of failures.
+We visualize the results in Figure 6.
+The naive approach without CCG and without TR failed a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed.
+The amount of failures dropped to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention.
+Finally, with both CCG and TR, we were able to successfully collect 128 episodes without any human intervention.
 IV-B Noise Augmentation Ablation
@@
 We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [40].
-We also qualitatively observe that policies trained with noise-augmented datasets tend to be better at re-adjusting the object with respect to the goal.
+We also qualitatively observed that policies trained with noise-augmented datasets were better at re-adjusting the object with respect to the goal.
 IV-C Comparison to Kinesthetic Teaching
-Figure 7: Comparison of success rates of policies trained using a dataset collected by LPP and a traditional kinesthetic teaching approach.
-We test models trained on an increasing amount of demonstration data.
-Policies trained on data collected by LPP outperform those trained on data collected from kinesthetic teaching in almost all dataset sizes.
+Figure 7: Comparison of success rates of policies trained using a dataset collected by PvP and a traditional kinesthetic teaching approach.
+We tested models trained on an increasing amount of demonstration data.
+Policies trained on data collected by PvP outperform those trained on data collected from kinesthetic teaching for almost all dataset sizes.
 The shaded region consists of one standard deviation.
-We compare the performance of policies trained using demonstration data from LPP and from kinesthetic teaching.
-In particular, we investigate whether the quality of the demonstrations differ.
-To compare, we train two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with LPP and another collected with kinesthetic teaching [39].
-We use a Gaussian mixture model as the action representation of both policies.
-For LPP, we use the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
+We compared the performance of policies trained using demonstration data from PvP and from kinesthetic teaching.
+In particular, we investigated whether the quality of the demonstrations differ.
+To compare, we trained two separate policies, both sharing the architecture outlined in Section III-E, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [41].
+We used a Gaussian mixture model as the action representation of both policies.
+For PvP, we used the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
 […]
-We hypothesize that the kinesthetic policy performs more poorly than the LPP policy because the overall quality of the human demonstrations is lower than the programmatic or machine-generated demonstrations produced by LPP.
+We hypothesize that the kinesthetic policy performed more poorly than the PvP policy because the overall quality of the human demonstrations was lower than the programmatic or machine-generated demonstrations produced by PvP.
 Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [42, 43, 39].
-In particular, human demonstrations can be sub-optimal due to human mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think).
+In particular, human demonstrations can be sub-optimal due to mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think).
+The authors of [39] note that the relative average trajectory length is a good proxy for the quality of the dataset.
+The average number of time steps for each demonstration was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset.
+The numbers in the brackets denote variations of one standard deviation.
+Kinesthetic demonstrations were on average longer and had greater variance in length.
+In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-D, which a human demonstrator cannot emulate consistently.
 V Limitations and Future Work
-The quality and diversity of the collected grasps in LPP are partly determined by the viewpoint of the camera.
+The quantity, quality and diversity of the collected grasps in PvP are partly determined by the viewpoint of the camera.
 In this work, we empirically select a viewpoint that is adequate for our use-case.
-It would be interesting to allow the manipulator to explore and find vantage points for better grasp generation.
-Furthermore, while tactile sensing is used to improved the robustness of data collection, it could also be used as an extra modality for the policy.
+It would be interesting to allow the manipulator to explore and find better grasp poses.
+Furthermore, while tactile sensing is used to improve the robustness of data collection, it could also be used as an extra modality for the policy.
+We have introduced a framework for data collection and tested it with two real-world tasks.
+It would be interesting to test PvP on a larger set of tasks.
+Lastly, PvP is a language-driven data collection method, and it would be natural to integrate it into a high-level language based planner [9].
 VI Conclusion
-We presented LPP, a self-supervised data collection method to collect expert demonstrations for placement tasks.
-LPP leverages recent advancement in grasp planners and the symmetry between picking and placing to autonomously collect placing demonstrations.
-LPP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing.
-We showed that the quality of the demonstrations produced by LPP surpasses that of human demonstrations collected by kinesthetic teaching.
-LPP provides a promising and simple direction for real world robotic placing demonstration data collection with minimal human labour.
+We presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks.
+PvP leverages recent advancements in grasp planners, tactile sensing, compliant control and the symmetry between picking and placing to autonomously collect placing demonstrations.
+PvP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing.
+We showed that the quality of the demonstrations produced by PvP surpasses that of human demonstrations collected by kinesthetic teaching.
+PvP provides a promising and pragmatic direction for real world robotic placing demonstration data collection with minimal human labour.

### v4

@@
 Abstract
-We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific contact-constrained locations.
+We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific, contact-constrained locations.
 With PvP, we approach the collection of robotic object placement demonstrations by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems.
 Specifically, we obtain placing demonstrations from a set of grasp sequences of objects initially located at their target placement locations.
-Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention by combining two modules: tactile regrasping and compliant control for grasps.
-We train a policy directly from visual observations through behavioral cloning, using the autonomously-collected demonstrations.
+Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention using two modules: compliant control for grasping and tactile regrasping.
+We train a policy directly from visual observations through behavioural cloning, using the autonomously-collected demonstrations.
 By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table).
-We validate our approach in home robotic scenarios that include dishwasher loading and table setting.
-Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
+We validate our approach in home robot scenarios that include dishwasher loading and table setting.
+Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of success rate and data efficiency, while requiring no human supervision.
 I Introduction
-Imitation learning (IL) provides an appealing, simple, and practical option by learning end-to-end control policies from expert demonstrations [1, 2].
-The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics.
-While there have been many attempts to transfer this paradigm to robotic systems [3, 4], robots have yet to achieve similar successes.
-A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect the large amount of expert demonstrations in the physical world.
+Imitation learning (IL) provides a scaleable, simple, and practical option to learn control policies from expert demonstrations [1, 2].
+The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics [3, 4].
+There have been many attempts to transfer this paradigm to IL for robotic systems [5, 6], but similar successes have yet to be achieved.
+A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect a large number of expert demonstrations in the physical world.
 Having said this, many recent methods attempt to streamline [7, 8] or automate [9, 10, 11, 12] data collection.
@@
 In this work, we are interested in investigating autonomous data collection for robotic object placement.
-As opposed to prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints.
-We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place if handled properly.
+As opposed to much prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints.
+We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place, if handled properly.
 We propose a novel approach named placing via picking (PvP) that automates the collection of expert demonstrations for a large subset of contact-constrained placing problems.
 We do so by taking advantage of a powerful grasp planner [13], tactile sensing, and compliant control.
-PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems.
-Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory.
+PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place tasks.
+Given an environment with objects initially at their target locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory.
 We generate demonstration data for object placement in a self-supervised manner—the picking phase provides the training supervision for the placing task.
 While appearing deceptively easy at first, we highlight the crucial importance of two modules, compliant control for grasping (CCG) and tactile regrasping (TR).
-CCG and TR enables (1) robust and uninterrupted pick-and-place and (2) leveraging environment contact-constraints as guides to inform object placement.
+CCG and TR enable (1) robust and uninterrupted pick-and-place and (2) the use of environment contact-constraints as guides to inform object placement.
 LABEL:fig:title provides a visual summary of our approach.
-Our main contributions are:
+Our main contributions are
   1. 1.
@@
   2. 2.
-    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection without human intervention in contact-constrained environments;
+    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection in contact-constrained environments;
   3. 3.
@@
 II Related Work
-In this section, we review previous approaches to automate robot data collection in the real world and other work that generally leverage the idea of “working backwards”.
+In this section, we review prior work on automated robot data collection, discuss the general idea of “working backwards,” and investigate existing robotic placement strategies.
 II-A Automatic Data Collection
@@
 Examples include using trajectory optimization [14, 15] and, closer to our work, simple scripted pick-and-place policies [9, 10, 11] as forms of supervision.
-The authors of [7] use a scripted pick-and-place policy to bootstrap the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator.
-In [9], the authors present AutoRT, a method that uses a combination of scripted, learned and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data.
+The authors of [9] use a scripted pick-and-place policy to bootstrap the collection of the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator.
+In [11], the authors present AutoRT, a method that uses a combination of scripted, learned, and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data.
 Using a combination of a large language model (LLM) and visual language models (VLMs), AutoRT observes scenes in the wild and comes up with plausible tasks.
-Once a task is defined and chosen, AutoRT samples a data collection policy to attempt the task and collect real world data.
-In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with potential contact constraints.
-Unlike prior work, we do not restrict the placing task to simpler small objects on flat horizontal surfaces.
-The authors of [9] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; Our work is complementary with many of the prior frameworks and PvP could be used as a prior policy within them.
+Once a task is defined and chosen, AutoRT samples a data collection policy to attempt to gather real world data for the task.
+In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with contact constraints.
+Unlike prior work, we do not restrict the placing task to simple small objects on flat horizontal surfaces.
+The authors of [11] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; our work is complementary to many of the previous frameworks, where PvP could be used as a prior policy.
+Other works have explored automatic data generation by bootstrapping from a single human demonstration [16, 17].
+In [16], a robot manipulation task is modelled as having two phases: a coarse approach trajectory phase towards a bottleneck pose, followed by a fine interaction trajectory phase.
+The single human demonstration provides the relative pose of the manipulator with respect to the task-relevant object, from which a self-supervised data collection procedure is formulated to train a bottleneck pose predictor.
+In our work, we also automatically collect and augment approach trajectory data.
+However, we do not require any human demonstrations since we focus on placing tasks only.
+By doing so, we can leverage the inherent symmetry of picking and placing for self-supervised data collection through the use of a combination of tactile sensing, compliant control, and an off-the-shelf grasp planner.
 II-B Working Backwards
-The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [14] and in learning visual predictive models for control [15, 16, 17].
+The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18], learning visual predictive models for control [19, 20, 21], and reinforcement learning [22].
 Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25].
 These methods consider objects that are assumed to be in their desired place configuration during data collection.
-The manipulator can then pick an object and move back to the initial goal location to generate data.
-However, we focus on placement as opposed to insertion as our task.
-Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple, generally larger objects and place poses involved in the task.
+The manipulator can then pick an object and move back to the initial goal location to generate training data.
+However, we focus on placement as opposed to insertion.
+Insertion policies generally only reason about local interactions.
+On the other hand, placement involves reasoning over a larger scene with the potential for multiple place poses and more complex objects that require planners to grasp successfully.
 Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning.
 Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner.
-However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement.
+However, we focus on generating demonstrations for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement.
 In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
 II-C Robotic Manipulator Placing
-Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [22, 23, 24], while placing is relatively less explored.
+Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [27, 28, 29], while placing has been relatively less studied.
 Previous approaches have explored the use of motion planning combined with place-specific objectives [30, 31].
@@
 The place problem has also been studied from the perspective of geometric modelling [32], where the planar surfaces of a discretized object model are matched to planar patches in the environment.
-Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time.
-Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [28] and imitation learning (IL) [29].
-In [29], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing.
-The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with any methods that require expert demonstrations.
+Importantly, 3D models of the objects being placed as well as the environment are assumed to be available ahead of time.
+Closer to our approach, other works have trained end-to-end place policies using reinforcement learning [33] and imitation learning [34].
+In [34], the authors demonstrate an inverse reinforcement learning technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing.
+The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with methods that require expert demonstrations.
 III Placing via Picking
-We first present an overview of our self-supervised data collection method, PvP, in Section III-A and then describe a method for noise augmentation during data collection in Section III-D.
-Next, we provide details about the procedure used to train our IL policy in Section III-E.
+We first present an overview of our self-supervised data collection method, PvP, in Section III-A, and then describe a method for noise augmentation during data collection in Section III-B.
+Finally, we provide details about the procedure used to train our IL policy in Section III-C.
 III-A Self-Supervised Data Collection
-[…]
-[…]
-[…]
-[…]
+(a) Grasp planning.
+(b) Grasping.
+(c) Retrieving.
+(d) Placing by reversing.
 Figure 1: The four steps involved in PvP, our autonomous demonstration data collection process for placing.
-We (a) generate grasps with an off-the-shelf grasp planner [11], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
-PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Figure 1.
+We (a) generate grasps with an off-the-shelf grasp planner [13]; (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp via tactile sensing; (c) retrieve the object with rotational compliance while storing the trajectory; and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
+PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Fig.
+1.
 During training, the objects of interest are initially in their goal positions.
 III-A1 Grasp Planning
-[…]
-[…]
-[…]
-[…]
+(a) Raw image.
+(b) Text-based object detection.
+(c) Segmentation.
+(d) Pruned grasps.
 Figure 2: The steps involved in our language-driven grasp planning pipeline.
-We use (a) Grounding-Dino [30] for object bounding box detection based on text and (b) Segment Anything [31] on the cropped images to produce object specific masks.
-We then use (c) Contact-graspnet [11] for grasp generation on only the segmented objects (i.e., masked areas).
+We use (b) Grounding-Dino [35] for object bounding box detection based on text descriptions and (c) Segment Anything [36] on the cropped images to produce object specific masks.
+We then use (d) Contact-GraspNet [13] for grasp generation on only the segmented objects (i.e., masked areas).
 Top: using “green plate” and “blue plate” as the targets for data collection.
 Bottom: using “bowl” and “mug” as the targets for data collection.
-At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1.
-We use the grasp planner Contact-GraspNet [11] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$.
-Given a set of text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [30] to find the object bounding boxes.
-Finally, we run Segment Anything [31] on each of the previously generated image crops from the bounding boxes to generate per-object masks.
-Contact-GraspNet can then use these masks to filter grasps that are not on objects of interest and keep only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$.
-We visualize an example of this process in Figure 2 for two different scenes.
+At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Fig.
+1(a).
+We use the grasp planner Contact-GraspNet [13] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in\mathrm{SE}(3)$ for all $i\in\{0,...,L\}$.
+Given a set of text-based descriptions of the objects of interest (e.g., “plates,” “green plate,” “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [35] to find the object bounding boxes.
+Finally, we run Segment Anything [36] on each of the previously-generated image crops from the bounding boxes to generate per-object masks.
+Contact-GraspNet then uses these masks to filter grasps that are not on objects of interest and keeps only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$.
+We visualize an example of this process in Fig.
+2 for two different scenes.
 After pruning, we randomly select a single grasp pose, $\mathbf{T}_{\text{grasp}}$, from $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$.
-To approach the object of interest without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
+To approach without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
 III-A2 Grasping
 Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory.
-An example of a grasp sequence is shown in Figure 1.
+An example of a grasp sequence is shown in Fig.
+1(b).
 During grasping, two modules play a critical role in the robustness of PvP: compliant control for grasping (CCG) and tactile regrasping (TR).
-III-B Compliant Control for Grasping
 We use a Cartesian impedance controller for manipulation.
@@
 Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values.
-The low compliance of the robot arm then allows the manipulator and gripper to comply with the contact constraints and minimize its applied forces against the environment.
+We call this CCG.
+The high compliance of the robot arm allows the manipulator and gripper to comply with the contact constraints and minimize the applied forces against the environment.
 Larger contact-constrained objects generate unnecessary reaction forces when gripped without compliance.
-These large forces trigger emergency stops and can damage the environment, and the robot.
-Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper ultimately leading to inaccurate placements.
-On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator and gripper towards a natural object placement pose.
-Figure 3: Tactile images from before and after a tactile regrasp (TR).
-Left: the plate’s contact surface area fills half of the tactile image indicating a shallow grasp.
-Right: after a regrasp, the contact surface area has increased indicating a stable grasp.
-[…]
-[…]
-[…]
-[…]
-Figure 4: Visualization of contact surfaces in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
-III-C Tactile Regrasping
+Large forces, in turn, trigger emergency stops and can damage the environment and the robot.
+Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper, ultimately leading to inaccurate placements.
+On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator towards a natural object placement pose.
 Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip significantly once grasped.
-In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) results in grasps that are unstable (e.g., a shallow grasp) and not ideal for PvP.
-To mitigate this issue, we use tactile sensing (specifically, a visuotactile Finger-STS [32] sensor) to preemptively detect if a grasp is stable and perform a regrasp if needed.
+In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) can result in grasps that are unstable (e.g., shallow grasps) and not ideal for PvP.
+To mitigate this issue, we use tactile sensing—specifically, a visuotactile Finger-STS [37] sensor—to preemptively detect if a grasp is stable and perform a regrasp if needed.
 The Finger-STS is capable of providing multimodal feedback (i.e., both tactile and visual observations).
-To this end, we design our regrasping module to detect the contact surface area and calculate the relative change in EE pose required to recover a larger contact surface and, thereby, a more stable grasp.
+To this end, we design our TR module to detect the contact surface area and calculate the relative change in end effector (EE) pose required to recover a larger contact surface and, thereby, a more stable grasp.
 We detect regions of contact based on marker displacements and the RGB image from the Finger-STS.
-If an impartial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose.
+If partial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose.
 We note that a simpler and more general random regrasp strategy could also be used.
-An example of this is shown in Figure 3.
-We visualize contact surfaces of various household items in Figure 4.
-In Section IV-A, we experimentally investigate the effects of including these modules through ablation studies.
-III-C1 Retrieving
+An example of this is shown in Fig.
+3.
+We visualize contact regions for various household items in Fig.
+4.
+In Section IV-A, we experimentally investigate the effects of including these modules through an ablation study.
+III-A3 Retrieving
 We begin the retrieval phase when a valid, stable grasp is detected.
-An image of the retrieval procedure is shown in Figure 1.
-We store the poses of the end effector (EE) as well as the respective timestamps throughout the retrieval process.
+An image of the retrieval procedure is shown in Fig.
+1(c).
+We store the poses of the EE as well as the respective timestamps throughout the retrieval process.
 For our work, we measure joint encoder readings at a rate of 120 $\mathrm{H}\mathrm{z}$.
 The manipulator is set to be compliant along the rotational axes of the EE only, when commanded to move to the pregrasp pose.
-This allows the manipulator to better comply with any unexpected environment interactions.
-Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ centered around a set fixed pose above the scene.
+Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ around a set fixed pose above the scene.
 That is, given a set fixed pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$ to recover a clearance pose.
-We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection.
-At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
+We use $\sigma_{tr}$ = 2.5 cm during data collection.
+At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose,
 $\tau^{\text{expert}}_{r}=((\mathbf{T}_{0},t_{0}),...,(\mathbf{T}_{M},t_{M})),$ (1)
@@
 To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a sparser trajectory $\bar{\tau}^{\text{expert}}_{r}$.
-We extract a total of $\bar{M}+1$ states where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$.
-We do this by finding the nearest pose in terms of time from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
+We extract a total of $\bar{M}+1$ states, where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$, by finding the nearest pose (in terms of time) from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
 $\bar{\tau}^{\text{expert}}_{r}=((\bar{\mathbf{T}}_{0},0),(\bar{\mathbf{T}}_{1},\Delta t),...,(\bar{\mathbf{T}}_{\bar{M}},\bar{M}\Delta t)).$ (2)
-In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
-III-C2 Placing by reversing
-We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory
+In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 s.
+III-A4 Placing by Reversing
+We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory,
 $\bar{\tau}^{\text{expert}}_{p}=(\bar{\mathbf{T}}_{\bar{M}},...,\bar{\mathbf{T}}_{0}).$ (3)
-An example of a place trajectory is shown in Figure 1. Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
-$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}_{\bar{M}-i-1},$ (4)
-where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in SE(3)$ is a relative change in EE pose.
+An example of a place trajectory is shown in Fig. 1(d). Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
+$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}^{\phantom{-1}}_{\bar{M}-i-1},$ (4)
+where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in\mathrm{SE}(3)$ is a relative change in EE pose.
 We define the expert action as $\mathbf{a}_{i}=(\Delta\mathbf{T}_{i},a_{\text{gripper},i})$, where $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close).
@@
 We downsample the images to a resolution of 128 $\times$ 128 and also stack the three previous RGB frames, which results in $\mathbf{o}_{i}\in\mathbb{R}^{128\times 128\times 12}$.
-During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$).
+During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$).
 Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper (i.e.
@@
 with a total length of $T=\bar{M}+N$.
-III-D Noise-Augmented Data Collection
-To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise.
-We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [33], and actions [34].
-We also draw inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [35], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment.
+Figure 3: Tactile images from before and after a tactile regrasp (TR).
+Left: the contact surface area of the plate fills half of the tactile image, indicating a shallow grasp.
+Right: after a regrasp, the contact surface area has increased, indicating a stable grasp.
+Figure 4: Visualization of contact surface region in the tactile image of grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
+III-B Noise-Augmented Data Collection
+To increase the coverage of our demonstrations and train more robust policies, we perturb the first 75% of the poses with noise.
+We take inspiration from previous works that have investigated the use of noise injection for both the policy observations [38], and actions [39].
+We also take inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [40], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment.
 We call this variant noise-augmented data collection.
@@
 $\mathbf{t}_{\text{perturbed}}=\mathbf{t}+\delta\mathbf{t},$ (6)
-where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation with
+where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is an isotropic Gaussian noise vector and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$.
+Similarly, we perturb the rotation with
 $\boldsymbol{\theta}_{\text{perturbed}}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (7)
-where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$.
+where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is an isotropic Gaussian noise vector and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$.
 We perturb both the rotation axis and the rotation angle.
-We find $\sigma_{t}=0.5~{}$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5~{}$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values.
+We find $\sigma_{t}=$ 0.5 cm, $\sigma_{e}=$ 0.5 cm and $\sigma_{\theta}=$ 0.5 to be reasonable values.
 We evaluate the effects of this augmentation on policy performance in Section IV-B
-III-E Policy Learning
-We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP.
-We use behavioral cloning (BC) [36] to train our policy with a likelihood-based loss or objective function.
+III-C Policy Learning
+We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\mid\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP.
+We use behavioural cloning (BC) [41] to train our policy with a likelihood-based loss or objective function.
 Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
-$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})}.$ (8)
+$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}\mid\mathbf{o}_{t})}.$ (8)
 The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [42], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions.
-In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians.
-In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [38] and [39], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value of $61e-$ 6.
+In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}\!\!\mid\!\!\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians.
+In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [43] and [44], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value.
 Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss.
 IV Experiments
-[…]
-[…]
+(a) Dishrack placement task, which consists of placing multiple plates of varying physical properties in evenly spaced slots.
+(b) Table placement task, which consists of placing a bowl on a plate and a cup on a coaster.
 Figure 5: Sequence of images from roll outs of place policies trained using data collected with PvP.
 The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
-We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 5.
-In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle).
-In this section, we present our quantitative results for the more challenging dishrack loading task as it better represents the type of task, with additional contact constraints from the environment, that PvP was designed for.
-We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
+We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Fig.
+5.
+In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object, and the state of the current scene to decide where and how to place (e.g., place angle).
+In this section, we only present our quantitative results for the more challenging dishrack loading task as it better represents the type of task (with added contact constraints from the environment) that PvP was designed for.
+We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B, and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
 IV-A PvP Data Collection Robustness
-Figure 6: Ablation study on compliant control for grasping (CCG) and tactile regrasping (TR).
-We measured the average number of episodes that could be collected autonomously based on the number of failures.
-Both CCG and TR play a crucial role in reducing the number of failures to zero, which allowed us to collect the necessary amount of data seamlessly.
 We require a robust data collection loop to collect a large number of place demonstrations autonomously with PvP.
-We studied the effects of compliant control for grasping (CCG) and tactile regrasping (TR), as introduced in Section III, on the robustness of PvP.
-We observed two main causes, initially discussed in Section III-A2, that led to misplaced plates during data collection.
-First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack.
+We studied the effects of CCG and TR, as introduced in Section III, on the robustness of PvP.
+We observed two main causes of misplaced objects during data collection, initially discussed in Section III-A2, First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the environment.
 During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released.
-This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot.
+This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the environment.
 With PvP, we mitigate this failure mode with CCG.
@@
 We conducted an ablation study on CCG and TR where we ran self-supervised data collection for a total of 128 episodes and recorded the number of failures.
-We visualize the results in Figure 6.
+We visualize the results in Fig.
+6.
 The naive approach without CCG and without TR failed a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed.
-The amount of failures dropped to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention.
+The number of failures dropped to three with the addition of CCG, which raises the average to 43 episodes collected before each intervention.
 Finally, with both CCG and TR, we were able to successfully collect 128 episodes without any human intervention.
+Figure 6: Ablation study on compliant control for grasping (CCG) and tactile regrasping (TR).
+We measured the average number of episodes that could be collected autonomously based on the number of failures.
+Both CCG and TR play a crucial role in reducing the number of failures to zero, which allowed us to collect the necessary amount of data seamlessly.
 IV-B Noise Augmentation Ablation
-TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation using two different policy action representations.
+TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation and using two different policy action representations.
 The noise-augmented data collection procedure improves the performance of both deterministic and Gaussian mixture policies.
@@
 Yes 87.78(3.93) 62/70
-[…]
-We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [40].
+In Section III-B, we introduced a noise-augmented data collection variant of PvP.
+We investigated its contribution to the success rate of the policy.
+We collected two different datasets each containing 128 demonstrations, with and without added noise augmentation.
+We tested the effect of noise-augmented data on policies trained using two different action representations: a deterministic policy and a Gaussian mixture policy with five modes.
+We present the results in Table I.
+For each row, we averaged the success rate over three models trained from scratch with three different seeds.
+We used 20 or 30 rollouts to calculate the success rate of each model.
+The numbers in the brackets denote variation of one standard deviation.
+We also report the total successes summed over the three models.
+For both action representations, we found that training on noise-augmented data improved the performance of the final policy significantly.
+The deterministic policy had an improvement of approximately 10% while the mixture policy had a larger improvement of 15%.
+We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the actions allows the policy to better capture this by not having to average over multiple possible expert actions [45].
 We also qualitatively observed that policies trained with noise-augmented datasets were better at re-adjusting the object with respect to the goal.
@@
 In particular, we investigated whether the quality of the demonstrations differ.
-To compare, we trained two separate policies, both sharing the architecture outlined in Section III-E, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [41].
+To compare, we trained two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [46].
 We used a Gaussian mixture model as the action representation of both policies.
 For PvP, we used the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
-[…]
+In Fig.
+7, we visualize the success rates of both models trained with varying numbers of demonstrations.
+As done previously, we averaged the performance over three models with different seeds for each datapoint.
+We tested each model using eight rollouts.
+The models trained with data from PvP, collected without any human involvement, significantly outperform the models trained using kinesthetic data by about 20% for almost all numbers of demonstrations.
+We observed that policies trained with kinesthetic teaching often misplaced the plate (e.g., by not having the right place angle or not resting the plate in a stable manner with respect to the environment contact points) and struggled to open the gripper at the correct times.
+The latter of the two error modes did not appear in policies trained with PvP.
 We hypothesize that the kinesthetic policy performed more poorly than the PvP policy because the overall quality of the human demonstrations was lower than the programmatic or machine-generated demonstrations produced by PvP.
@@
 The authors of [44] note that the relative average trajectory length is a good proxy for the quality of the dataset.
-The average number of time steps for each demonstration was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset.
-The numbers in the brackets denote variations of one standard deviation.
+The average number of time steps for all demonstrations was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset.
+The numbers in the brackets denote variation of one standard deviation.
 Kinesthetic demonstrations were on average longer and had greater variance in length.
-In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-D, which a human demonstrator cannot emulate consistently.
+In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-B, which a human demonstrator cannot emulate consistently.
 V Limitations and Future Work
-The quantity, quality and diversity of the collected grasps in PvP are partly determined by the viewpoint of the camera.
-In this work, we empirically select a viewpoint that is adequate for our use-case.
+PvP relies on effective grasping (i.e., accurate grasp detections and stable grasps during retrieval).
+This makes this method susceptible to the typical challenges involved in grasping, namely dealing with noisy point cloud measurements, unreliable object segmentation, and unmodelled object dynamics (e.g.
+object slip).
+The quantity, quality, and diversity of the grasp poses found by the planner are partly determined by the viewpoint of the camera.
+In this work, we empirically selected a viewpoint that is adequate for our use case.
 It would be interesting to allow the manipulator to explore and find better grasp poses.
@@
 We have introduced a framework for data collection and tested it with two real-world tasks.
-It would be interesting to test PvP on a larger set of tasks.
+We would like to test PvP on a larger set of tasks for a more exhaustive evaluation.
 Lastly, PvP is a language-driven data collection method, and it would be natural to integrate it into a high-level language based planner [11].
 VI Conclusion
-We presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks.
-PvP leverages recent advancements in grasp planners, tactile sensing, compliant control and the symmetry between picking and placing to autonomously collect placing demonstrations.
+In this paper, we presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks.
+PvP leverages recent advancements in grasp planners, tactile sensing, compliant control, and the symmetry between picking and placing to autonomously collect placing demonstrations.
 PvP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing.
 We showed that the quality of the demonstrations produced by PvP surpasses that of human demonstrations collected by kinesthetic teaching.
-PvP provides a promising and pragmatic direction for real world robotic placing demonstration data collection with minimal human labour.
+PvP provides a promising and pragmatic direction to collect real world robotic placing demonstration data with minimal human labour.